how to Find the slope and y-intercept of the line whose equation is given.
4(y - 3) + (x - 7) = 5(x + 1) - 5

Found 2 solutions by Alan3354, Boreal:
Answer by Alan3354(69443)

how to Find the slope and y-intercept of the line whose equation is given.
4(y - 3) + (x - 7) = 5(x + 1) - 5
------------
Solve for y.
That'll give you y = mx + b
m is the slope, b is the y-intercept

Answer by Boreal(15235)

4(y - 3) + (x - 7) = 5(x + 1) - 5
put it in the form of y=mx+b
4y-12+x-7=5x+5-5
4y-19=4x
4y=4x+19
y=x+(19/4); slope is 1; y-intercept is 19/4
check
when x=0, y=(19/4)
19-12-7=0=0 to check